how do you write this problem in slope intercept form y+5=3/2(x-3)
1
Expert's answer
2014-06-02T07:13:55-0400
The slope intercept form:

y=mx+b. (1)

Firstly, multiply 3/2 by (x-3): y+5=3x/2-9/2.

Add to both sides of the equation -5: y=3x/2-19/2.
We obtain that m=3/2, b= -19/2.

Answer. y=3x/2-19/2.